OpenStudy (owlstudies):

In the following problem, what are the domain restrictions? **i worked out the answer but I'm not sure if 5 would be a restriction or not since it was in the denominator in some steps of the problem**

OpenStudy (owlstudies):

\[\frac{ 2\frac{ x+3 }{ x-5 }+7 }{ \frac{ x+3 }{ x-5 }-6 }\]

OpenStudy (owlstudies):

I simplified it to : \[\frac{ 9x-29 }{ -5x+33 }\]

OpenStudy (owlstudies):

One of the domain restrictions is 33/5, but would the other one be 5, or no?

OpenStudy (mathmate):

Yes, 5 is also a restriction, because it occurs in the denominator of the numerator (sounds confusing, but you understand what I mean). If there were no restriction, one would be dividing a zero by a zero.

OpenStudy (owlstudies):

okay thank you!

OpenStudy (mathmate):

You're welcome! :)